40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6474     Entry time: Sat Mar 31 08:01:07 2012
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: DRMI measurement 

I have measured the sensing matrix of the DRMI although the lock still doesn't stay for a long time.

As for the noise budget, it looks very tough as there are more glitches than that in the PRMI.

In this weekend I will take some more trials in the DRMI lock until I am satisfied.

ELOG V3.1.3-